OpenStudy (anonymous):

Find all values of x such that sin (2x) = sin (x) and 0 < x < 2\(\pi\)

ganeshie8 (ganeshie8):

sin2x = sinx 2sinxcosx = sinx sinx(2cosx-1) = 0

ganeshie8 (ganeshie8):

use zero product property

OpenStudy (anonymous):

hmm so you have sinx = 0 and cosx = 1/2

OpenStudy (anonymous):

can u get that from a graph?

OpenStudy (anonymous):

can you?

ganeshie8 (ganeshie8):

yes work it

OpenStudy (anonymous):

anyway...that's 2 roots down...how about the three others?

OpenStudy (anonymous):

i think it is the easier way since sin(2x) have period of pi u can graph both graph and see the common points

OpenStudy (anonymous):

graphing sinusoidal functions? seems tedious

OpenStudy (anonymous):

@ganeshie8 do you know how to get the three other roots of this?

OpenStudy (anonymous):

hmm seems like he doesn't. he ran away =_= lol

ganeshie8 (ganeshie8):

sinx(2cosx-1) = 0 sinx = 0 ; cosx = 1/2 x = 0, pi ; x = pi/3

ganeshie8 (ganeshie8):

We're done. since you want to find roots within \((0, 2 \pi) \)

ganeshie8 (ganeshie8):

btw 0 wont work. so you would have oly 2 roots x = pi, pi/3

OpenStudy (anonymous):

 this is the way that i know the graphs meet in 0,pi and 2pi but i dont know why i got a different answer ro @ganeshie8